40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1955     Entry time: Thu Aug 27 12:34:48 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: up to arm power 70 
Last night, I tried to run locking scripts.
The power went up to 70 a couple of times .
Then it failed to switch to RF CARM.
I was too tired at that time to figure out what is the problem with the switching.
But it seemed to me that the problem could be solved by some gain tweaking.
Looks like the IFO is back to a good state.
ELOG V3.1.3-